OpenStudy (anonymous):

40 dm(cube) of H gass was collected over water at 831 torr pressure at 23 degree centi.What would be the volume of dru hydrogen gas at standard conditions?The vapour pressure of water at 23 degree centi. is 21 torr.

OpenStudy (anonymous):

vol. of moist \(H_2\)=40 \(dm^3\) P=831 Torr T=\(23degree\)\(centigrade\) P\(H_2O\) at 23\(degree~centigrade\)=21 torr vol. of dry \(H_2\)=? \(\large{solution;}\) how we can solve it by finding \(\Huge{v_1}\)=\(\Huge?\ \Huge{v_2}\)=\(\Huge?\) and others @aaronq @thomaster have a look here please

OpenStudy (aaronq):

So you would subtract the partial pressure of the water vapour from the total pressure to get the pressure due to \(H_{2~(g)}\) only. From there you can use: \(\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}\)

OpenStudy (anonymous):

but how we find \(\P_1=?~P_2=?~~T_1~and~T_2=?\) please help me more

OpenStudy (aaronq):

\(P_2\) and \(T_2\) are given in the question (i.e standard conditions) \(P_1=P_{H_2}=P_{T}-P_{H_2O}\)

OpenStudy (anonymous):

i am solving tring

OpenStudy (aaronq):

degrees centigrade means degrees Celsius. We need to convert Celsius/centigrade to Kelvin Kelvin = degrees celsius + 273\(^oC\)

OpenStudy (anonymous):

\(\Huge\color{purple}{Thank~You}\) @aaronq can you please solve it more above very useful information you tell me i mean about kelvin

OpenStudy (aaronq):

So, 23\(^oC\) to Kelvin. 23\(^oC\) + 273\(^oC\) = 296 Kelvin

OpenStudy (anonymous):

yes.... next .....solve it completely by understanding me...!

OpenStudy (aaronq):

okay, so \(\sf P_1=P_{H2}=P_T-P_{H_2O}=831~ torr-21~torr=811~torr\) \(\dfrac{811~torr*40~L}{296~K}=\dfrac{P_2V_2}{T_2}\)

OpenStudy (aaronq):

plug in \(P_2\), \(T_2\) and solve

OpenStudy (anonymous):

once again thankyou i solve it@aaronq

OpenStudy (aaronq):

no problem!

OpenStudy (anonymous):

@aaronq @aaronq @aaronq

OpenStudy (anonymous):

how u learn chemistery tell me this trick @aaronq i learn one line 200 times but very difficulty i learn and physics as well .......can you please help me to get rid of this problem ....and give me any suggestion?

OpenStudy (aaronq):

It's about knowing the concepts then practising them. You get better with every question you complete.

OpenStudy (anonymous):

very easy questions like Rutherford model and others that i understan...but when i learn after few days ...it becomes out from my mind ...very easy questions out from my mind ...y its happend with me..~i don't know @aaronq

OpenStudy (aaronq):

yeah, if you dont keep seeing it, you're very likely to forget it. Thats part of the reason why I come on this site.